If $8,000.00 is deposited in a bank that compounds interest anually at 5%, then what is the interest paid at the end of the second year?
Answer by stanbon(57347)

If $8,000.00 is deposited in a bank that compounds interest anually at 5%, then what is the interest paid at the end of the second year?
--------
A(t) = P(1+(r/n))^(nt)
A(2) = 8000(1+(0.05)/2))^(1*2)
A(2) = 8000(1+0.025)^2
A(2) = $8405.00
----------------------
Amount of interest earned is $405.00
==============================================
Cheers,
Stan H.